# Shapes and orientations of massive halos in the statistically anisotropic universe

### v1

@@ III.1 The halo shapes
$\mathcal{G}_{ij}\equiv g_{*}\left(\hat{d}_{i}\hat{d}_{j}-\frac{1}{3}\delta_{ij}\right).$ (5)
@@ III.2 The halo orientations
Figure 4: PDFs of $\hat{A}_{x},~{}\hat{A}_{y}$ and $\hat{A}_{z}$, i.e., the halo orientations, for the halos with masses of $14<\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\leq 14.5$ from the L2 run. Each color line shows the result for $g_{*}=0$ (blue), ±$\pm 0.1$ (green), ±$\pm 0.5$ (red) and ±$\pm 1.0$ (orange), and the dashed (solid) lines are for the cases with the positive (negative) sign.
@@
As we have discussed in the previous subsection, we found that the SA effect on the halo shape is negligible. To explore whether there are any halo properties that are affected by the SA, we next examine the orientation of the halo as an ellipsoid. In Fig. 4, we show the PDFs of the components of the vector $\mathbf{\hat{A}}$, which characterizes the orientation of the largest axis of each halo, and here we use the halos with masses in the range of $14<\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\leq 14.5$, taken from the L2 run only. Each color line shows the result for $g_{*}=0$ (blue), ±$\pm 0.1$ (green), ±$\pm 0.5$ (red) and ±$\pm 1.0$ (orange), and the dashed (solid) lines are for the cases with the positive (negative) sign. Note that $\hat{A}_{x}$ and $\hat{A}_{y}$ take values in the range $[-1,~{}1]$, while $\hat{A}_{z}$ takes values in the range $[0,~{}1]$ due to our setup described in Eq. (3).
@@
In the isotropic case, i.e., $g_{*}=0$, $P(\hat{A}_{x}),~{}P(\hat{A}_{y})$ and $P(\hat{A}_{z})$ are completely flat, indicating that the distribution of the orientations of halos is uniformly random. [^2] This does not imply that the spatial correlations of the halo orientations are zero, as discussed in the context of the intrinsic alignment of galaxies [32, 33, 34].. Unlike the shape parameters shown in Figs. 2 and 3, the halo orientations exhibit a clear dependence on $g_{*}$. As $g_{*}$ increases positively, both $P(\hat{A}_{x})$ and $P(\hat{A}_{y})$ exhibit increasing curvature, developing deeper minima at $\hat{A}_{x}=0$ and $\hat{A}_{y}=0$, and more pronounced maxima at $\hat{A}_{x}=\pm 1$ and $\hat{A}_{y}=\pm 1$, respectively. In contrast, $P(\hat{A}_{z})$ shows the opposite trend, peaking at $\hat{A}_{z}=0$ and reaching a minimum at $\hat{A}_{z}=1$. These behaviors indicate that, for larger positive values of $g_{*}$, halos as ellipsoids tend to align parallel to either the $x$-axis or the $y$-axis.
Conversely, as $g_{*}$ decreases (i.e., becomes more negative), $P(\hat{A}_{z})$ becomes a steeper increasing function of $\hat{A}_{z}$, exhibiting a higher peak at $\hat{A}_{z}=1$ and a lower minimum at $\hat{A}_{z}=0$. Correspondingly, $P(\hat{A}_{x})$ and $P(\hat{A}_{y})$ show the opposite behavior, peaking at $\hat{A}_{x}=0$ and $\hat{A}_{y}=0$ and reaching minima at $\hat{A}_{x}=\pm 1$ and $\hat{A}_{y}=\pm 1$, respectively. This means that, for negative values of $g_{*}$, halos tend to align parallel to the $z$-axis.
@@
In summary, halo orientations tend to increasingly align perpendicular to $\hat{\mathbf{d}}$ for larger positive values of $g_{*}$, and parallel to $\hat{\mathbf{d}}$ for lower negative values. This trend is consistent with the visualizations of matter distributions on large scales presented in Fig. 1. As discussed in the previous subsection, the SA in our setup can be considered as the effect of the global tensor field, and then the alignment of the halos can be affected by the global structure (see, e.g., Ref. [16]).
@@
Let us mention a few behaviors of PDFs observed in Fig. 4, that are related to the symmetry in our setup. First, $P(\hat{A}_{x})$ and $P(\hat{A}_{y})$ are nearly identical. This is because the $x$- and $y$-axes are both perpendicular to the SA direction $\hat{\mathbf{d}}=(0,~{}0,~{}1)$, and there is no statistical distinction between the $x$- and $y$-directions. Second, $P(\hat{A}_{x}),~{}P(\hat{A}_{y})$, and $P(\hat{A}_{z})$ are not symmetric with respect to the sign of $g_{*}$, e.g., between $g_{*}=+1$ and $g_{*}=-1$. As discussed above, in the case of $\hat{\mathbf{d}}=(0,~{}0,~{}1)$, halos align in parallel to $x$- or $y$-axes for positive values of $g_{*}$ while $z$-axis for negative values. Consequently, the maximum amplitudes of $|P(\hat{A}_{x})-0.5|,~{}|P(\hat{A}_{y})-0.5|$ and $|P(\hat{A}_{z})-1|$ are not equal between positive and negative $g_{*}$; they are suppressed for positive $g_{*}$ because the halo orientations are split between two perpendicular axes.
@@
Figure 5: PDFs of $\hat{A}_{x}$, $P(\hat{A}_{x})$, for the halo mass ranges of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~{}13.5],~{}(14,~{}14.5]$ and $(15,~{}15.5]$. Each line shows the results for $g_{*}=0$ and ±$\pm 0.1$.
Figure 6: Same as Fig. 5 but for the PDFs of $\hat{A}_{z}$, $P(\hat{A}_{z})$.
@@
Figs. 5 and 6 show the mass dependence of the modification of PDFs of $\hat{A}_{x}$ and $\hat{A}_{z}$ due to the SA, respectively. From top to bottom, we show the results for the halo mass ranges $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~{}13.5]$$(14,~{}14.5]$, and $(15,~{}15.5]$. The correspondence between line type and $g_{*}$ values is the same as in Fig. 4. We omit $P(\hat{A}_{y})$ here, as it is nearly identical to $P(\hat{A}_{x})$ due to the symmetry. For $g_{*}=\pm 0.1$, $P(\hat{A}_{x})$ and $P(\hat{A}_{z})$ show negligible deviations from the isotropic case in the lowest mass range, while the deviations become more pronounced in the $(14,~{}14.5]$ bin and even stronger in $(15,~{}15.5]$, maintaining the same $g_{*}$-dependence as in Fig. 4. These results suggest stronger SA-induced alignment for more massive halos. This trend is physically reasonable, as more massive halos have larger volumes roughly proportional to their mass and can more effectively “feel” global effects such as SA.
Figure 7: PDF of $\hat{A}_{z}$ for the high- and low-$c_{\rm vir}$ halos of $14<\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\leq 14.5$ from the L2 run with $g_{*}=\pm 0.1$.
@@
We also examine whether additional factors influence the alignment of halos induced by the SA. In particular, we investigate the correlation between the halo formation epoch and its alignment. It is well established that the halo concentration parameter serves as a reliable proxy for the formation epoch [35]. We define the halo concentration parameter as
$c_{\rm vir}=R_{\rm vir}/r_{\rm s}~{},$ (6)
@@
where $R_{\rm vir}$ and $r_{\rm s}$ are, respectively, the virial radius of the halo and the scale radius of the NFW density profile, $\rho_{\rm NFW}(r)=\rho_{\rm s}/[(r/r_{\rm s})(1+r/r_{\rm s})^{2}]$ [36]. We adopt the values of $R_{\rm vir}$ and $r_{\rm s}$ obtained from the Rockstar halo finder. We divide the halos of $14<\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\leq 14.5$ from the L2 run with $g_{*}=\pm 0.1$ into two groups based on $c_{\rm vir}$: those with concentrations larger and smaller than the median value ${\tilde{c}}_{\rm vir}$. Fig. 7 shows the distribution function of $\hat{A}_{z}$ for the halos with high and low $c_{\rm vir}$. For both values of $g_{*}$, halos with higher and lower concentrations exhibit similar alignment distributions, indicating that the formation epoch does not significantly affect halo alignment associated with the SA. We found the same trend when halos were divided based on the triaxiality parameter $T$ as well as for halos with masses in the range in $15<\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\leq 15.5$ in the $g_{*}=\pm 0.1$ case. These results suggest that the initial conditions influenced by SA are the primary factor responsible for the halo alignment in the statistically anisotropic universe.
IV conclusion
@@
In this work, we investigated the impact of the SA in the matter density field on the shapes and orientations of cluster-sized halos at $z=0$ using cosmological $N$-body simulations that incorporate a quadrupolar-type SA characterized by the magnitude parameter $g_{*}$. For halo shapes, we examined the three-dimensional axis ratio $s$ (smallest-to-largest) and the triaxiality parameter $T$. Halo orientations were represented by the direction of the largest axis as ellipsoid, denoted by the vector $\mathbf{\hat{A}}$. We found that, while both of the shape parameters, $s$ and $T$, show little sensitivity to SA (Figs. 2 and 3), the halo orientations are significantly affected (Fig. 4). We also found that halo orientations tend to align perpendicular to the SA-preferred direction $\hat{\mathbf{d}}$ for positive values of $g_{*}$, and parallel for negative values (Fig. 4).
@@
We used halos with masses of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~{}13.5]$$(14,~{}14.5]$, and $(15,~{}15.5]$, and investigated the mass dependence of the alignment due to the SA. We clearly showed that this SA alignment effect becomes more pronounced for more massive halos, suggesting that more massive halos are more sensitive to the global anisotropic nature of the matter distribution (Figs. 5 and 6). Importantly, this alignment is expected to originate from the initial conditions that have inherent information about the SA, rather than later-stage nonlinear processes, as we found no significant dependence on halo properties such as concentration or ellipticity (Fig. 7).
@@
In this paper, we employed a very simple scale-independent quadrupole type of SA, but other types of SA are actually possile. In fact, in the context of the inflation models with gauge fields including the anisotorpic inflation, the possibility of generating various types of SA has been actively discussed, e.g., statistically anisotropic non-Gaussianity (see, e.g., Ref. [37]), the SA in the primordial gravitational waves (see, e.g., Ref. [38]), and so on. In conjunction with those observations, it would be interesting to apply our analysis to more specific model-based types of SA to verify the early universe.
@@
As the PDFs presented in this paper are one-point functions, they are insensitive to spatial correlations in halo orientations. However, the SA-induced changes in halo orientations may have implications for other cosmological observables, such as intrinsic alignments (IA) [32, 33, 34] (see also Refs. [39, 40] for recent studies on the IA of galaxy clusters). Since these effects could be potential sources of systematic errors in cosmological analyses or serve as tools to constrain the SA parameter, investigating how SA influences a broader range of statistical measures is an important direction for future work.
@@
Our findings highlight the potential of using halo alignment as a complementary probe of the SA in the Universe. In particular, measurements of projected halo ellipticity, which is observable in galaxy cluster-galaxy lensing observations, would exhibit systematic directional features due to the SA-induced alignment and could be used to constrain $g_{*}$. Further works, including projection effects, redshift evolution, and connections to observable quantities, will be necessary to assess the feasibility of such constraints in future surveys.
Acknowledgements.

### v2

@@ 3.1 Halo shapes
$\mathcal{G}_{ij}=g_{*}\left(\hat{d}_{i}\hat{d}_{j}-\frac{1}{3}\delta_{ij}\right).$ (3.2)
@@ 3.2 Halo orientations
Figure 3: PDFs of $\hat{A}_{x},~\hat{A}_{y}$ and $\hat{A}_{z}$, i.e., the halo orientations, for the halos with masses of $14<\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\leq 14.5$ from the L2 run. Each color line shows the result for $g_{*}=0$ (blue), $\pm 0.1$ (green), $\pm 0.5$ (red), and $\pm 1.0$ (orange), and the dashed (solid) lines are for the cases with the positive (negative) sign.
@@
As we discussed in the previous subsection, we found that the effect of SA on the halo shape is negligible. To explore whether there are any halo properties that are affected by the SA, we next examine the orientation of the halo as an ellipsoid. In Fig. 3, we show the PDFs of the components of the vector $\mathbf{\hat{A}}$, which characterizes the orientation of the largest axis of each halo. Here we use the halos with masses in the range of $14<\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\leq 14.5$, taken from the L2 run only. Each color line shows the result for $g_{*}=0$ (blue), $\pm 0.1$ (green), $\pm 0.5$ (red), and $\pm 1.0$ (orange), and the dashed (solid) lines are for the cases with the positive (negative) sign. Note that $\hat{A}_{x}$ and $\hat{A}_{y}$ take values in the range [1,1][-1,~1], while $\hat{A}_{z}$ takes values in the range [0,1][0,~1] due to our setup described in Eq. (2.3).
@@
In the isotropic case, i.e., $g_{*}=0$, $P(\hat{A}_{x}),~P(\hat{A}_{y})$ and $P(\hat{A}_{z})$ are completely flat, indicating that the distribution of the orientations of halos is uniformly random. [^2] This does not imply that the spatial correlations of the halo orientations are zero, as discussed in the context of the intrinsic alignment of galaxies [croft00, heavens00, lee00].. Unlike the shape parameters shown in Fig. 2, the halo orientations exhibit a clear dependence on $g_{*}$. As $g_{*}$ increases positively, both $P(\hat{A}_{x})$ and $P(\hat{A}_{y})$ exhibit an increasing curvature, developing deeper minima at $\hat{A}_{x}=0$ and $\hat{A}_{y}=0$, and more pronounced maxima at $\hat{A}_{x}=\pm 1$ and $\hat{A}_{y}=\pm 1$, respectively. In contrast, $P(\hat{A}_{z})$ shows the opposite trend, peaking at $\hat{A}_{z}=0$ and reaching a minimum at $\hat{A}_{z}=1$. These behaviors indicate that, for larger positive values of $g_{*}$, halos as ellipsoids tend to align parallel to either the $x$-axis or the $y$-axis. As $g_{*}$ decreases (i.e., becomes more negative), $P(\hat{A}_{z})$ becomes a steeper increasing function of $\hat{A}_{z}$, showing a higher peak at $\hat{A}_{z}=1$ and a lower minimum at $\hat{A}_{z}=0$. Similarly, $P(\hat{A}_{x})$ and $P(\hat{A}_{y})$ show the opposite behavior, peaking at $\hat{A}_{x}=0$ and $\hat{A}_{y}=0$ and reaching minima at $\hat{A}_{x}=\pm 1$ and $\hat{A}_{y}=\pm 1$, respectively. This means that, for negative values of $g_{*}$, halos tend to align parallel to the $z$-axis.
@@
In summary, halo orientations tend to increasingly align perpendicular to $\hat{\mathbf{d}}$ for larger positive values of $g_{*}$, and parallel to $\hat{\mathbf{d}}$ for lower negative values. This trend is consistent with the visualizations of matter distributions on large scales presented in Fig. 1. As discussed in the previous subsection, the SA in our setup can be considered as the effect of the global tensor field, and then the alignment of the halos can be affected by the global structure (see, e.g., Ref. [shiraishi23]).
@@
Let us mention a few behaviors of the PDFs observed in Fig. 3 that are related to the symmetry in our setup. First, $P(\hat{A}_{x})$ and $P(\hat{A}_{y})$ are nearly identical. This is because the $x$- and $y$-axes are both perpendicular to the SA direction $\hat{\mathbf{d}}=(0,~0,~1)$, and there is no statistical distinction between the $x$- and $y$-directions. Second, $P(\hat{A}_{x}),~P(\hat{A}_{y})$, and $P(\hat{A}_{z})$ are not symmetric with respect to the sign of $g_{*}$, e.g., between $g_{*}=+1$ and $g_{*}=-1$. As discussed above, in the case of $\hat{\mathbf{d}}=(0,~0,~1)$, halos align in parallel to the $x$- or $y$-axes for positive values of $g_{*}$ while the $z$-axis for negative values. Consequently, the maximum amplitudes of $|P(\hat{A}_{x})-0.5|,~|P(\hat{A}_{y})-0.5|$ and $|P(\hat{A}_{z})-1|$ are not equal between positive and negative $g_{*}$; they are suppressed for positive $g_{*}$ because the halo orientations are split between two perpendicular axes.
@@
Figure 4: Left: PDFs of $\hat{A}_{x}$, $P(\hat{A}_{x})$, for the halo mass ranges of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~13.5],~(14,~14.5]$ and (15,15.5](15,~15.5]. Each line shows the results for $g_{*}=0$ and $\pm 0.1$. Right: Same as the left panel but for the PDFs of $\hat{A}_{z}$, $P(\hat{A}_{z})$.
@@
Fig. 4 shows the mass dependence of the modification of the PDFs of $\hat{A}_{x}$ and $\hat{A}_{z}$ due to the SA, respectively. From top to bottom, we show the results for the halo mass ranges $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~13.5]$(14,14.5](14,~14.5], and (15,15.5](15,~15.5]. The correspondence between line type and $g_{*}$ values is the same as in Fig. 3. We omit $P(\hat{A}_{y})$ here, as it is nearly identical to $P(\hat{A}_{x})$ due to the symmetry with respect to $\hat{\mathbf{d}}$. For $g_{*}=\pm 0.1$ which are comparable to the upper limit currently obtained from the galaxy clusteing measurements [sugiyama18], $P(\hat{A}_{x})$ and $P(\hat{A}_{z})$ show negligible deviations from the isotropic case in the lowest mass range, while the deviations become more pronounced in the (14,14.5](14,~14.5] bin and even stronger in (15,15.5](15,~15.5], maintaining the same $g_{*}$-dependence as in Fig. 3. The errors are larger in the halo mass range of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(15,~15.5]$ because the number of halos for this mass range is approximately ten times smaller than that for $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(14,~14.5]$. These results suggest a stronger SA-induced alignment for more massive halos. This trend is physically reasonable, as more massive halos have larger volumes roughly proportional to their mass and can more effectively “feel” global effects such as SA.
Figure 5: PDFs of $\hat{A}_{z}$ for the high- and low-$c_{\rm vir}$ halos of $14<\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\leq 14.5$ from the L2 run with $g_{*}=\pm 0.1$.
@@
We also examine whether additional factors influence the alignment of halos induced by the SA. In particular, we investigate the correlation between the halo formation epoch and its alignment. It is well established that the halo concentration parameter serves as a reliable proxy for the formation epoch [Wechsler02]. We define the halo concentration parameter as
$c_{\rm vir}=R_{\rm vir}/r_{\rm s}~,$ (3.3)
@@
where $R_{\rm vir}$ and $r_{\rm s}$ are, respectively, the virial radius of the halo and the scale radius of the NFW density profile, $\rho_{\rm NFW}(r)=\rho_{\rm s}/[(r/r_{\rm s})(1+r/r_{\rm s})^{2}]$ [NFW]. We adopt the values of $R_{\rm vir}$ and $r_{\rm s}$ obtained from the Rockstar halo finder. We divide the halos of $14<\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\leq 14.5$ from the L2 run with $g_{*}=\pm 0.1$ into two groups based on $c_{\rm vir}$: those with concentrations larger and smaller than the median value ${\tilde{c}}_{\rm vir}$. Fig. 5 shows the distribution function of $\hat{A}_{z}$ for the halos with high and low $c_{\rm vir}$. For both values of $g_{*}$, halos with higher and lower concentrations exhibit similar alignment distributions, indicating that the formation epoch does not significantly affect the halo alignment associated with the SA. We found the same trend when the halos were divided according to the triaxiality parameter $T$ as well as for halos with masses in the range $15<\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\leq 15.5$ in the $g_{*}=\pm 0.1$ case. These results imply that the halo alignment mainly reflects the SA-imprinted initial conditions, rather than the subsequent formation process.
3.3 Other vectors associated with the dynamics of halos
In this section, we extend our investigation to additional vector quantities associated with the dynamics of halos: bulk velocities and angular momenta, in order to examine whether there exist vector quantities with SA-dependent directionality besides the orientation of the halo as an ellipsoid. We are also motivated to investigate these quantities because SA-distorted matter distributions can affect the matter and halo bulk velocity fields.
3.3.1 Halo bulk velocities
Figure 6: PDFs of the halo bulk velocities $\hat{V}_{x}$ (left top), $\hat{V}_{y}$ (right top), and $\hat{V}_{z}$ (left bottom), for the halos with masses of $14<\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\leq 14.5$ from the L2 run. Each color line shows the result for $g_{*}=0$ (blue), $\pm 0.1$ (green), $\pm 0.5$ (red) and $\pm 1.0$ (orange). The dashed (solid) lines stand for the positive (negative) sign.
To examine whether the halo bulk velocity vector, ${\bf{V}}=(V_{x},~V_{y},~V_{z})$, carries any imprint of SA, we focus on the unit vector of the bulk velocity of a halo $\hat{{\bf V}}={\bf V}/|{\bf V}|$, and study the PDFs of each component for several values of $g_{*}$. Fig. 6 shows the PDFs of the bulk velocity component within the mass range $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(14,~14.5]$. In this figure, we observe a similar trend to that seen in the PDFs of halo orientations in Fig. 3, indicating that the bulk velocity of the halo is also affected by the SA. Note that the dependence on the sign of $g_{*}$ of the bulk velocity vector distributions is observed to be opposite to that of the orientations. In the case of $g_{*}<0$ ($g_{*}>0$), structures form preferentially parallel (perpendicular) to the $z$-direction, and the associated gravitational potential wells are also algined along the parallel (perpendicular) directions to the $z$-axis. Consequently, halos are more likely to fall into these potential wells, which increases the number of bulk-velocity vectors aligned perpendicular (parallel) to the $z$-axis. However, it is clear that sensitivity to the SA of the bulk velocity vectors is still weaker compared to the orientations. This is more clearly shown in Fig. 7. Fig. 7 illustrates how the effect of SA on the bulk velocity depends on halo mass for $g_{*}=\pm 0.1$. We find that the bulk velocity vectors do not significantly depend on the SA with respect to the size of the statistical error. We note that the probability distributions $P({\hat{V}_{x}})$ for the mass range of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~13.5]$, and $\in(14,~14.5]$ exhibit a mildly downward convex profile around $\hat{V}_{x}=0.0$ even for the case with $g_{*}=0$. Further data and realizations will be required to confirm this feature with higher statistical significance.
Figure 7: Left: PDFs of $\hat{V}_{x}$, $P(\hat{V}_{x})$, for the halo mass ranges of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~13.5],~(14,~14.5]$ and (15,15.5](15,~15.5]. Each line is the result for $g_{*}=0$ and $\pm 0.1$. Right: Same as the left panel but for the PDFs of $\hat{V}_{z}$, $P(\hat{V}_{z})$.
3.3.2 Halo angular momenta
We also study the angular momentum vector of each halo provided in the Rockstar halo catalog, ${\bf{J}}=(J_{x},~J_{y},~J_{z})$ in the SA universe. To highlight the directional information of the angular momentum without being dominated by its amplitude, we focus on the unit vector of the angular momenta: $\hat{\bf J}={\bf J}/|{\bf J}|$. In Fig. 8, we present the distributions of the components of $\hat{\bf J}$ for halos in the mass range $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(14,~14.5]$. Each panel includes results for various values of the SA parameter $g_{*}=0,~\pm 0.1,~\pm 0.5,~\pm 1.0$, allowing us to visually inspect potential trends. In this figure, a similar trend to that seen in the PDFs of halo bulk velocities in Fig. 6 can be observed, indicating that the direction of the angular momentum of the halo is also affected by the SA. Note that the dependence on the sign of $g_{*}$ of the angular momentum distributions is observed to be opposite to that of the orientations. This behavior originates from the trend for the angular momentum of halos to be oriented parallel to its minor-axis vector as already found in Refs. [1992ApJ...399..405W, Bailin:2004wu, Shaw:2005dy, Bett:2006zy, 2006MNRAS.367.1781A], though our results demonstrate that this characteristic persists even in the presence of SA. However, it is clear that sensitivity to the SA of the angular momentum is much weaker compared to the orientations and bulk velocities. Fig. 9 shows the halo-mass dependence of the impact of SA on the halo angular momenta $\hat{J}_{x}$ and $\hat{J}_{z}$ for $g_{*}=\pm 0.1$. The comparison across mass ranges reveals that there is no statistically significant dependence of the SA effect on the angular momentum for the $g_{*}$ that is comparable to the upper limit from the observations of galaxy clustering [sugiyama18]. Consequently, the angular momentum vector is a less sensitive than the orientation and bulk velocity for probing the SA effect.
When taking into account previous sections and Fig. 4 (orientations) together with the results from Fig. 7 (bulk velocities) and Fig. 9 (angular momenta), it becomes evident that orientation is the most sensitive probe of SA among the quantities in this work, examined for $g_{*}=\pm 0.1$ which are comparable to the upper limit currently obtained from the galaxy clusteing measurements [sugiyama18].
Figure 8: PDFs of $\hat{J}_{x},~\hat{J}_{y}$ and $\hat{J}_{z}$, i.e., the halo angular momenta vector, for the halos with masses of $14<\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\leq 14.5$ from the L2 run. Each color line shows the result for $g_{*}=0$ (blue), $\pm 0.1$ (green), $\pm 0.5$ (red), and $\pm 1.0$ (orange), and the dashed (solid) lines are for the cases with the positive (negative) sign.
Figure 9: Left: PDFs of $\hat{J}_{x}$, $P(\hat{J}_{x})$, for the halo mass ranges of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~13.5],~(14,~14.5]$ and (15,15.5](15,~15.5]. Each line is the result for $g_{*}=0$ and $\pm 0.1$. Right: Same as the left panel but for the PDFs of $\hat{J}_{z}$, $P(\hat{J}_{z})$.
4 Conclusion
@@
In this work, we investigated the impact of the SA in the matter density field on the shapes and orientations of cluster-sized halos at $z=0$ using cosmological $N$-body simulations that incorporate a quadrupolar-type SA characterized by the magnitude parameter $g_{*}$. For halo shapes, we examined the three-dimensional axis ratio $s$ (smallest-to-largest) and the triaxiality parameter $T$. Halo orientations were represented by the direction of the largest axis as an ellipsoid, denoted by the vector $\mathbf{\hat{A}}$. We found that, while both of the shape parameters, $s$ and $T$, show little sensitivity to SA (Fig. 2), the halo orientations are significantly affected (Fig. 3). We also found that halo orientations tend to align perpendicular to the SA-preferred direction $\hat{\mathbf{d}}$ for positive values of $g_{*}$, and parallel for negative values (Fig. 3).
@@
We used halos with masses of $\log_{10}[M_{\rm 200b}/(h^{-1}M_{\odot})]\in(13,~13.5]$(14,14.5](14,~14.5], and (15,15.5](15,~15.5], and investigated the mass dependence of the alignment due to the SA. We clearly showed that this SA alignment effect becomes more pronounced for more massive halos, suggesting that more massive halos are more sensitive to the global anisotropic nature of the matter distribution (Fig. 4). Importantly, this alignment is expected to originate from the initial conditions that have inherent information about the SA, rather than later-stage nonlinear processes, as we found no significant dependence on halo properties such as concentration or ellipticity (Fig. 5). We also investigated the sensitivity of other vector quantities related with the dynamics of halos such as the halo bulk velocity and angular momentum, to the SA effect. Although their PDFs show qualitatively similar responses to the SA, which is opposite to that of the halo orientations, the magnitude of these changes is smaller than that seen in the halo orientations. This comparison leads us to conclude that the orientations are the most sensitive among the examined quantities in this paper, and thus serve as a promising new probe of SA.
@@
In this paper, we employed a simple scale-independent quadrupole type of SA, but other types of SA are actually possible. In fact, in the context of the inflation models with gauge fields including the anisotropic inflation, the possibility of generating various types of SA has been actively discussed, e.g., statistically anisotropic non-Gaussianity (see, e.g., Ref. [Yokoyama:2008xw]), the SA in the primordial gravitational waves (see, e.g., Ref. [Fujita:2018zbr]), and so on. In conjunction with those observations, it would be interesting to apply our analysis to more specific model-based types of SA to verify the early universe.
@@
As the PDFs presented in this paper are one-point functions, they are insensitive to spatial correlations in halo orientations. However, the SA-induced changes in halo orientations may have implications for other cosmological observables, such as intrinsic alignments (IA) [croft00, heavens00, lee00] (see also Refs. [shi24, ishikawa25] for recent studies on the IA of galaxy clusters). Since these effects could be potential sources of systematic errors in cosmological analyses or serve as tools to constrain the SA parameter, investigating how SA influences a broader range of statistical measures is an important direction for future work.
@@
Our findings highlight the potential of using halo alignment as a complementary probe of the SA in the Universe. In particular, measurements of projected halo ellipticity, which is observable in galaxy cluster-galaxy lensing observations, would exhibit systematic directional features due to the SA-induced alignment and could be used to constrain $g_{*}$. Further issues, including projection effects, redshift evolution, and connections to observable quantities, will be necessary to assess the feasibility of such constraints in future surveys.
Acknowledgements